OpenStudy (anonymous):

The height a ball bounces is less than the height of the previous bounce due to friction. The heights of the bounces form a geometric sequence. Suppose a ball is dropped from one meter and rebounds to 95% of the height of the previous bounce. What is the total distance travel by the ball when it comes to rest? *Does the problem give you enough information to solve the problem? *How can you write the general term of the sequence? *What formula should you use to calculate the total distance?

OpenStudy (anonymous):

@Data_LG2

OpenStudy (zzr0ck3r):

so the series is \(1+0.95(1) + 0.95^2+0.95^3.......)

OpenStudy (anonymous):

okay what would the general term of the sequence?

OpenStudy (zzr0ck3r):

so \(\sum_0^\infty (0.95)^n\)

OpenStudy (anonymous):

okay what is the total distance then??

OpenStudy (zzr0ck3r):

The sum of a geometric series with ratio less than one is given by \(\frac{a}{1-r}\) where \(a\) is the first term, and \(r\) is the common ratio.

OpenStudy (zzr0ck3r):

wait it would be 1+2(.95)+2(.95)(.95).....

OpenStudy (anonymous):

2*.95^1+2*.95^2+...+2*.95^inifinity which is = to -(2*.95)/(.95-1)=38 + 1 :P which is the initial drop =39

OpenStudy (anonymous):

right?

OpenStudy (zzr0ck3r):

it would be \(1+2\sum_1^\infty (0.95)^n\)

OpenStudy (zzr0ck3r):

I get 39

OpenStudy (zzr0ck3r):

good job